# Towards Global Optimal Visual In-Context Learning Prompt Selection

### v1

@@ 4.3 Model analysis
Is quality of in-context examples really aligned with visual similarity? A basic claim in VPR is that the more similar an image is to the query image, the better it is as an in-context example. Nonetheless, VPR utilizes performance metrics instead of visual similarity for contrastive learning and provides better results, which makes one wonder if such a claim is indeed solid. To check this out, we take a look at the best performing in-context examples selected by our method. The results are shown in Fig. 4(a) and (b). We first visualize the correlation between visual similarity, which is computed with CLIP features, and mIoU on segmentation for both VPR and our method. It is clear that visual similarity can be a basic heuristic for choosing visual in-context examples. Generally well performing examples appear to share high similarity with query samples. On the other hand, there are also large amount of failure cases with high similarity, which indicates the ground principle should be much more complex than visual similarity. Based on this conclusion, we further visualize several cases in Fig. 4(c) selected by our method, which better shows that one should consider more than visual similarity and stress other factors such as object size, spatial position. On the opposite, we find that the background similarity, which can contribute significantly to the visual similarity, is hardly related to the quality of in-context examples.

### v2

@@ 4.3 Model analysis
Is quality of in-context examples really aligned with visual similarity? A basic claim in VPR is that the more similar an image is to the query image, the better it is as an in-context example. Nonetheless, VPR utilizes performance metrics instead of visual similarity for contrastive learning and provides better results, which makes one wonder if such a claim is indeed solid. To check this out, we take a look at the best performing in-context examples selected by our method. The results are shown in Fig. 4(a) and (b). We first visualize the correlation between visual similarity, which is computed by extracting CLIP features for each image and then calculating and averaging the cosine similarity between the alternatives and query, and mIoU on segmentation for both VPR and our method. It is clear that visual similarity can be a basic heuristic for choosing visual in-context examples. Generally well performing examples appear to share high similarity with query samples. On the other hand, there are also large amount of failure cases with high similarity, which indicates the ground principle should be much more complex than visual similarity. Based on this conclusion, we further visualize several cases in Fig. 4(c) selected by our method, which better shows that one should consider more than visual similarity and stress other factors such as object size, spatial position. On the opposite, we find that the background similarity, which can contribute significantly to the visual similarity, is hardly related to the quality of in-context examples.
@@ Appendix B Additional ablation study
Transferability of Partial2Global. To test the transferability of the prompt selection method and further reveal the potential of our method, we can add the demonstration suggested by conducting the following experiment for both SupPR and our method: we use models trained on each fold of segmentation and apply it to other folds. The results are shown in Tab. 6. We find that while both two methods degrade in the transfer learning setting, our method still outperforms SupPR in general. This interesting results indicate that, the training data size of each fold is insuifficient for training a robust and generalizable ranking model. On the other hand, this again indicates that prompt selection for VICL cannot be simply based on visual similarity, as claimed in our paper.
Table 5: Cross-fold performance of our method on segmentation task.
Source/Target Fold 0 1 2 3
0 36.38 32.63 30.90
1 35.74 32.94 31.32
2 34.16 36.16 30.44
3 34.28 35.93 32.98
Table 6: Cross-fold performance of SupPR on segmentation task.
Source/Target Fold 0 1 2 3
0 35.46 32.44 30.95
1 34.92 32.96 31.03
2 34.71 36.48 30.08
3 34.01 35.83 32.15
Efficiency of our method. One would ask if the proposed method would suffer from poor efficiency, compared with SupPR. In general, the usage of list-wise ranker and the ranking aggregation will inevitably introduce additional computational cost, while the increased complexity during inference is affordable under common circumstances. Specifically, we provide the training and inference time cost as follows. (1) The training of list-wise ranker on the colorization task, which contains about 500000 ranking sequences, takes about 10 hours on 8 V100s. Once the model is trained it can be directly used for any other queries with the same ranking criteria as the training task without any further finetuning. (2) During inference on one V100 gpu, our proposed pipeline requires about 1.17s to rank 50 alternatives for each query sample in the complete process, including feature extracting (0.3s), sub-sequence ranking with list-wise ranker (0.8s), and ranking aggregation (0.07s). Note that some techniques could be utilized to accelerate this process. For example, when we prepare the extracted features in advance (which is reasonable given the candidate set can be prepared in advance), we can skip the feature extraction stage and reduce the time cost by 0.3s. With engineering works, the inference time cost can be further reduced. The detailed inference speed (feature extraction included) given different alternative set size is presented as in Tab. 7.
Table 7: Inference speed with different alternative set size.
alternative set size inference time for each query (s)
25 1.03
50 1.17
100 1.40
Upper bound of selecting different in-context prompts. Another interesting problem is to examine the ’upper bound’ of performance by adopting different in-context prompts. To further provide insight to this task, we try to examine an upper bound: directly testing all alternatives for each query in the segmentation task and presenting the best IoU in Tab. 8. While our proposed method is much better than SupPR, it still leaves great potential for better performance, which we will take as future works.
Table 8: Oracle in-context learning performance on segmentation.
fold0 fold1 fold2 fold3
SupPR 37.08 38.43 34.40 32.32
Ours 38.81 41.54 37.25 36.01
best iou among 50 alternative (oracle) 48.75 52.62 49.75 49.03
Sensitivity test for hyper-parameters. In general, our method is robust against changes of hyper-parameters. To show this, we try the suggestions to conduct ablation studies on these two hyper-parameters, whose results are presented in the following table. For delta which denotes the margin, using 0 margin leads to worse results while larger margins would be better for learning ranking models. For tau which is the temperature coefficient in NeuralNDCG, we simply use the best setting tau=1 from the original paper. As can be seen in Tab. 9, using smaller temperature would not lead to better results. Nonetheless, all hyperparameter settings enjoy better performance than SupPR, validating the effectiveness of our method.
Table 9: Ablation study for hyper-parameters.
$\delta=0$ $\delta=2$ $\tau=0.01$ $\tau=0.1$
MSE 0.594 0.588 0.608 0.592
Impact of alternative set size on performance. Our choice of alternative set is the same as VPR, i.e. selecting 50 most visually similar samples based on CLIP features. Typically, this choice is good enough as shown in previous works like VPR. Here, to further investigate the impact of data quality, we thus tried the suggestion to test our method on all folds of segmentation task with 25 or 100 alternatives for each query. The results are shown in Tab. 10. One can find that our method is also very robust against different sizes of alternative set.
Table 10: Ablation study for different alternative set sizes.
set size fold0 fold1 fold2 fold3
25 38.48 41.82 37.14 35.60
50 (main) 38.81 41.54 37.25 36.01
100 38.81 41.80 37.90 36.00
Effectiveness of different terms in our proposed loss. We have conducted an additional ablation study to compare models trained for colorization without each loss term. The results are shown in Tab. 11. In general all three loss terms contribute the final performance, with $\mathcal{L}_{sort}$ plays the most important role.
Table 11: Ablation study for loss terms.
w/o $\mathcal{L}_{sort}$ w/o $\mathcal{L}_{margin}$ w/o $\mathcal{L}_{mse}$ full loss
MSE 0.601 0.595 0.585 0.583